# Learning to Generate Unit Tests for Automated Debugging

### v2

@@ Abstract
Unit tests (UTs) play an instrumental role in assessing code correctness as well as providing feedback to large language models (LLMs), motivating automated test generation. However, we uncover a trade-off between generating unit test inputs that reveal errors when given a faulty code and correctly predicting the unit test output without access to the gold solution. To address this trade-off, we propose UTGen, which teaches LLMs to generate unit test inputs that reveal errors along with their correct expected outputs based on task descriptions. Since model-generated tests can provide noisy signals (e.g., from incorrectly predicted outputs), we propose UTDebug that (i) scales UTGen via test-time compute to improve UT output prediction, and (ii) validates and backtracks edits based on multiple generated UTs to avoid overfitting, and helps LLMs debug effectively. We show that UTGen outperforms other LLM-based baselines by 7.59% based on a metric measuring the presence of both error-revealing UT inputs and correct UT outputs. When used with UTDebug, we find that feedback from UTGen’s unit tests improves pass@1 accuracy of Qwen2.5 32B on HumanEvalFix and our own harder debugging split of MBPP+ by over 3.17% and 12.35% (respectively) over other LLM-based UT generation baselines. Lastly, we demonstrate that UTGen is a better judge for code correctness, outperforming a state-of-the-art trained 8B reward model by 4.43% on HumanEval+ with best-of-10 sampling using Qwen2.5 7B.111Code: https://github.com/archiki/UTGenDebug
[ roundcorner=5pt, backgroundcolor=UserExampleBg, linecolor=UserExampleTitle, outerlinewidth=0.5pt, frametitlebackgroundcolor=UserExampleTitle, frametitlefont=, ]user_example
Learning to Generate Unit Tests for Automated Debugging
Archiki Prasad Elias Stengel-Eskin Justin Chih-Yao Chen Zaid Khan Mohit Bansal UNC Chapel Hill
@@ 1 Introduction
With rapid advancements in training large language models (LLMs; Achiam et al., 2023; Anthropic, 2024; Gemini et al., 2023), enhancing their coding abilities has garnered significant attention (Chen et al., 2021; Li et al., 2023; Roziere et al., 2023; Guo et al., 2024b, inter alia). However, these models are far from perfect and – much like human-written code – model-written code contains errors. Human developers often improve their code through test-driven development, i.e., identifying failure cases by providing example inputs and their expected outputs – referred to as unit tests (UTs) – to test individual functionalities, reasoning over causes of failure, and modifying the code to address the issues (Maximilien and Williams, 2003; Nagappan et al., 2008; Ficco et al., 2011; Beck, 2022). Models have similarly been shown to benefit from iterative debugging based on explicit or implicit feedback stemming from failing unit tests Chen et al. (2023b); Moon et al. (2023). However, the ability to provide feedback and debug incorrect code is often bottlenecked by the availability of (failing) unit tests for a given problem.
While several coding benchmarks come with human-written UTs for evaluation purposes, these suites of UTs are typically small due to the laborious annotation process Liu et al. (2024b). Unit test collection is challenging for two reasons: first, it requires sampling inputs that are likely to trigger an error. For example in Fig. 1, the unit test: (120)==121 would not trigger any error (despite the fact that the code is incorrect for non-multiples of 10), while another unit test: (123)==131 would lead to an incorrect output, revealing a bug in the function. Secondly, UTs require expected outputs (e.g., 131 in the previous UT), i.e., the desired behavior of the code being tested must be known. Due to these challenges, prior work employs LLMs to generate unit test inputs at random (i.e., without conditioning on faulty code to generate failing inputs) and often uses the gold (correct) code solution for generating outputs Chen et al. (2023a); Chae et al. (2024). Therefore, these approaches do not scale well at inference time, which requires an online and automated method for generating UTs as the LLMs solve problems on the fly. To bridge this gap, we pose the following research questions:
@@
To address RQ1, we characterize two desirable properties of unit test generators (in Sec. 3.1): 1) high attack rate, i.e., given a faulty code, the unit test generator should generate inputs that are likely to trigger errors;222Here, an ‘error’ means either a syntax/runtime error, or when the result of the target code does not match the expected output of the UT. We expand on this in Sec. 3. 2) high output accuracy, ensuring that the UT output is consistent with the task description (and that of a correct solution). For instance, in Fig. 1, (120) would lead to a lower attack rate, as it does not trigger any errors, while (123)==131 does (high attack rate); however, both have high output accuracy as in both cases the output is correct. We measure these properties via three intrinsic metrics: measuring attack rate and output accuracy independently and then, crucially, how often a UT generator can generate both failing inputs along with correct UT outputs based only on the problem description and a faulty code. We benchmark the ability of LLMs to act as zero-shot unit test generators and show that coding LLMs generally struggle (cf. Sec. 5.1).
@@
Moreover, addressing RQ2, we find that, without finetuning, zero-shot models often exhibit a strong trade-off between attack rate and output accuracy. In other words, the UTs that are most likely to trigger errors (i.e., higher attack rate) are generally more challenging edge cases, making it harder for the model to predict their output (i.e., lower output accuracy). Due to a lack of dedicated datasets for unit test generation in prior work, we introduce UTGen, a data creation and training recipe (Sec. 3.2) that bootstraps training data for UT generation from existing code generation datasets by perturbing code to simulate errors, generating failing unit test and augmenting it with chain-of-thought rationales Wei et al. (2022). We show that training via UTGen better balances this trade-off and yields a higher number of unit tests that have both attacking inputs and correct outputs, with 7.59% absolute improvement over prompting models to generate failing UTs.
@@
Finally, we examine RQ3, utilizing our generated UTs on downstream, i.e., extrinsic tasks: code debugging and code generation. Here, automated UT generation methods face additional challenges: unlike human-generated gold UTs – which have correct outputs but require human involvement – generated UTs provide noisy feedback, as the UT might fail to reveal the buggy code’s errors or have an incorrectly predicted output. To mitigate this issue, we propose UTDebug, an improved multi-turn debugging method that addresses these challenges in two ways (cf. Sec. 3.3): (i) we improve the output accuracy of generated UTs by scaling test-time compute via self-consistency (Wang et al., 2022); (ii) we regularize the debugging process by generating multiple UTs and accepting code edits only if the revised code passes more generated UTs, backtracking edits otherwise. We then plug UTs generated from multiple LLM-based methods into UTDebug in Sec. 5.2, finding that on our most challenging subset of MBPP+ problems, UTDebug with UTs generated by UTGen improves pass@1 accuracy of Qwen2.5 7B by 12.9% (absolute) compared to debugging without UTs, and by 5.3% over a zero-shot unit test generation baseline. We demonstrate that by generating UTs with UTGen and computing the rate of passing UTs, we can better judge code correctness than using scores from trained, state-of-the-art 8B reward model (RM). For best-of-10 sampling with Qwen2.5 7B, UTGen improves accuracy by 4.43% over the RM on HumanEval+.
@@ 2 Related Work
Automatic Unit Test Generation. Manually writing unit tests is laborious and often infeasible Chen et al. (2023a); Liu et al. (2024b). Consequently, past research explores automatic UT generation (King, 1976; Cadar et al., 2008; Holler et al., 2012; Cha et al., 2015, inter alia). The advent of LLMs has spurred recent efforts in using them for UT generation Chen et al. (2023a); Schäfer et al. (2023); Liu et al. (2024b). Specifically, Schäfer et al. (2023) and Liu et al. (2024b) focus on generating unit test inputs via prompting frontier LLMs like GPT-4 and/or iterative prompting, assuming access to the gold solution. In contrast, our models, trained with UTGen, generate both input-output UT pairs based on the task description without relying on the gold implementation. While Chen et al. (2023a) also generate input-output UT pairs using standard LLM prompting, their primary focus is code generation – not the quality of generated UTs. On the other hand, we directly model the desiderata of UTs including output accuracy, and demonstrate its utility on code generation and debugging.
LLM Debugging. Using LLMs for debugging faulty code, or program repair, has been extensively studied. Debugging approaches are divided into those training models to debug Moon et al. (2023); Ni et al. (2024); Chae et al. (2024) and those providing external feedback to pretrained models Chen et al. (2023b); Olausson et al. (2023); Zhong et al. (2024). Both rely on (gold) unit tests for training or feedback. Thus, UTGen complements both methods by providing generated unit tests when human-written tests are scarce or unavailable. In Sec. 3.3, we introduce UTDebug, a debugging pipeline that addresses noisy feedback from inaccurate unit tests through test-time scaling and backtracking. Moreover, in Sec. 5 we show that UTGen’s unit tests can effectively provide feedback to LLMs for code generation and debugging.
3 UTGen and UTDebug: Unit Test Generation and Automated Debugging
@@ 3.1 Background and Task Setup
Given a natural language task description for a coding task, we focus on generating unit tests (UTs) in the format of input-output pairs that are consistent with the task description . Our setup is consistent with Chen et al. (2023a) and Jain et al. (2024) who also consider unit tests in the form of input-output pairs generated without utilizing the correct implementation of the function. More generally, our setting is akin to parameterized unit testing Tillmann et al. (2010) and uses the notion of functional correctness, i.e., measuring correctness by simulating an exhaustive set of scenarios (UT inputs) and ensuring that the function performs as expected (as per the problem description ).
@@
3.2 Training LLMs for Unit Test Generation
While prior work focus on curating training data for improving code generation Guo et al. (2024a); Muennighoff et al. (2023), there is a general lack of dedicated datasets for training the desired UT generator outlined above. Therefore, to improve along the lines of RQ2, we design, UTGen, a training recipe that bootstraps this data from training datasets for code generation i.e., a collection of problem descriptions and their corresponding code implementations.
@@ 3.2 Training LLMs for Unit Test Generation
We start with a collection of coding problems with problem descriptions () and gold codes () as illustrated in Fig. 2 (I). Specifically, we use publicly available data from Tulu-3 Lambert et al. (2024a) due to its large scale and the improvements noted by Lambert et al. (2024a) when finetuning LLMs on coding tasks.333The code solutions in the Tulu-3 SFT mix have either been written by humans or by frontier models and are thus highly likely to be correct Lambert et al. (2024a). We filter it to focus on Python code with functional abstractions (further details in Appendix B). However, in order to train an adversarial unit test generator we would need access to incorrect or faulty code solutions that can be debugged, as unit tests must have some error to attack. We obtain these by using the LLM to perturb the reference code solution.
@@
As mentioned in Sec. 3.1, one of the goals of the unit test generator is to be able to not only generate valid unit tests, but failing unit tests (that trigger errors). To facilitate this, given a problem description, reference code , and buggy candidate code , we sample different unit test inputs (via the prompts in Appendix E). A unit test is failing if , i.e., if the output of the candidate fails to match the reference output (cf. Fig. 2 (II)). Note that, while the LLM can be used to generate the output of unit test, it can often be inaccurate Jain et al. (2024); Gu et al. (2024); so to ensure output accuracy, we use the output of the reference code during training. We revisit the discussion of how accurate LLMs are at output prediction during inference in Sec. 5.
Data Curation for Supervised Finetuning.
When training LLMs, the input to the LLM is the same prompt used for sampling unit tests (listed in Appendix E) and the output is an adversarial unit test. The goal here is to improve both output accuracy and attack rate jointly via supervised finetuning using the negative log-likelihood loss. Furthermore, we train the model to generate chain-of-thought (CoT; Wei et al., 2022) reasoning before predicting the UT output. To this end, we employ the post-hoc rationalization procedure outlined in Zelikman et al. (2022) – given the entire UT , we ask the LLM to generate rationales supporting why is the output corresponding to the input . Then, to create the supervision data, we add these rationales as CoTs prior to the output prediction (see Fig. 2 (III)).
3.3 Generated Unit Tests for Debugging
A key difference when using generated unit tests, as opposed to human-generated UTs, is the degree of noise in the feedback. Despite training models via UTGen or other methods, a generated unit test may not be 100% accurate. This can manifest in two ways: 1) the generated UT input does not fail for the code under debugging , i.e., ; 2) the generated UT output is inaccurate, i.e., not consistent with what a gold solution would yield (). Both types of errors can negatively impact the utility of unit tests for debugging, as shown in Fig. 3 (left). A non-adversarial input might result in faulty code being misclassified as correct and prematurely removed from debugging. Additionally, incorrect outputs can cause false positives, introducing errors to otherwise correct code. Even if the candidate code has bugs, incorrect UT outputs can lead to incorrect feedback, thereby, degrading performance. These issues motivate our RQ3 to incorporate noisy feedback. We propose UTDebug, which includes two ways to mitigate noisy feedback from automatically generated unit tests (detailed in Appendix C):

### v3

@@ Abstract
Unit tests (UTs) play an instrumental role in assessing code correctness as well as providing feedback to large language models (LLMs), motivating automated test generation. However, we uncover a trade-off between generating unit test inputs that reveal errors when given a faulty code and correctly predicting the unit test output without access to the gold solution. To address this trade-off, we propose UTGen, which teaches LLMs to generate unit test inputs that reveal errors along with their correct expected outputs based on task descriptions. Since model-generated tests can provide noisy signals (e.g., from incorrectly predicted outputs), we propose UTDebug that (i) scales UTGen via test-time compute to improve UT output prediction, and (ii) validates and backtracks edits based on multiple generated UTs to avoid overfitting, and helps LLMs debug effectively. We show that UTGen outperforms other LLM-based baselines by 7.59% based on a metric measuring the presence of both error-revealing UT inputs and correct UT outputs. When used with UTDebug, we find that feedback from UTGen’s unit tests improves pass@1 accuracy of Qwen2.5 32B on HumanEvalFix and our own harder debugging split of MBPP+ by over 3.17% and 12.35% (respectively) over other LLM-based UT generation baselines. Moreover, we observe that feedback from Qwen2.5 32B-based UTGen model can enhance debugging with frontier LLMs like GPT-4o by 13.8%. Lastly, we demonstrate that UTGen is a better judge for code correctness, outperforming a state-of-the-art trained 8B reward model by 4.43% on HumanEval+ with best-of-10 sampling using Qwen2.5 7B. Our code and datasets are publicly available at https://github.com/archiki/UTGenDebug.
@@ 1 Introduction
With rapid advancements in training large language models (LLMs; Achiam et al., 2023; Anthropic, 2024; Gemini et al., 2023), enhancing their coding abilities has garnered significant attention (Chen et al., 2021; Li et al., 2022; 2023; Roziere et al., 2023; Guo et al., 2024b, inter alia). However, these models are far from perfect and – much like human-written code – model-written code contains errors. Human developers often improve their code through test-driven development, i.e., identifying failure cases by providing example inputs and their expected outputs – referred to as unit tests (UTs) – to test individual functionalities, reasoning over causes of failure, and modifying the code to address the issues (Maximilien & Williams, 2003; Nagappan et al., 2008; Ficco et al., 2011; Beck, 2022). Models have similarly been shown to benefit from iterative debugging based on explicit or implicit feedback stemming from failing unit tests (Zhang et al., 2023; Chen et al., 2023b; Moon et al., 2023). However, the ability to provide feedback and debug incorrect code is often bottlenecked by the availability of (failing) unit tests for a given problem.
While several coding benchmarks come with human-written UTs for evaluation purposes, these suites of UTs are typically small due to the laborious annotation process (Liu et al., 2024b). Unit test collection is challenging for two reasons: first, it requires sampling inputs that are likely to trigger an error. For example in Fig. 1, the unit test: (120)==121 would not trigger any error (despite the fact that the code is incorrect for non-multiples of 10), while another unit test: (123)==131 would lead to an incorrect output, revealing a bug in the function. Secondly, UTs require expected outputs (e.g., 131 in the previous UT), i.e., the desired behavior of the code being tested must be known. Due to these challenges, prior work employs LLMs to generate unit test inputs at random (i.e., without conditioning on faulty code to generate failing inputs) and often uses the gold (correct) code solution for generating outputs (Chen et al., 2023a; Chae et al., 2024). Therefore, these approaches do not scale well at inference time, which requires an online and automated method for generating UTs as the LLMs solve problems on the fly. To bridge this gap, we pose the following research questions:
@@
To address RQ1, we characterize two desirable properties of unit test generators (in Sec. 3.1): 1) high attack rate, i.e., given a faulty code, the unit test generator should generate inputs that are likely to trigger errors;111Here, an ‘error’ means either a syntax/runtime error, or when the result of the target code does not match the expected output of the UT. We expand on this in Sec. 3. 2) high output accuracy, ensuring that the UT output is consistent with the task description (and that of a correct solution). For instance, in Fig. 1, (120) would lead to a lower attack rate, as it does not trigger any errors, while (123)==131 does (high attack rate); however, both have high output accuracy as in both cases the output is correct. We measure these properties via three intrinsic metrics: measuring attack rate and output accuracy independently and then, crucially, how often a UT generator can generate both failing inputs along with correct UT outputs based only on the problem description and a faulty code. We benchmark the ability of LLMs to act as zero-shot unit test generators and show that coding LLMs generally struggle (cf. Sec. 5.1).
@@
Moreover, addressing RQ2, we find that, without finetuning, zero-shot models often exhibit a strong trade-off between attack rate and output accuracy. In other words, the UTs that are most likely to trigger errors (i.e., higher attack rate) are generally more challenging edge cases, making it harder for the model to predict their output (i.e., lower output accuracy). Due to a lack of dedicated datasets for unit test generation in prior work, we introduce UTGen, a data creation and training recipe (Sec. 3.2) that bootstraps training data for UT generation from existing code generation datasets by perturbing code to simulate errors, generating failing unit test and augmenting it with chain-of-thought rationales (Wei et al., 2022). We show that training via UTGen better balances this trade-off and yields a higher number of unit tests that have both attacking inputs and correct outputs, with 7.59% absolute improvement over prompting models to generate failing UTs.
@@
Finally, we examine RQ3, utilizing our generated UTs on downstream, i.e., extrinsic tasks: code debugging and code generation. Here, automated UT generation methods face additional challenges: unlike human-generated gold UTs – which have correct outputs but require human involvement – generated UTs provide noisy feedback, as the UT might fail to reveal the buggy code’s errors or have an incorrectly predicted output. To mitigate this issue, we propose UTDebug, an improved multi-turn debugging method that addresses these challenges in two ways (cf. Sec. 3.3): (i) we improve the output accuracy of generated UTs by scaling test-time compute via self-consistency (Wang et al., 2022); (ii) we regularize the debugging process by generating multiple UTs and accepting code edits only if the revised code passes more generated UTs, backtracking edits otherwise. We then plug UTs generated from multiple LLM-based methods into UTDebug in Sec. 5.2, finding that on our most challenging subset of MBPP+ problems, UTDebug with UTs generated by UTGen improves pass@1 accuracy of Qwen2.5 Coder 32B Instruct by 15.07% (absolute) compared to debugging without UTs, and by 4.61% over a zero-shot unit test generation baseline (with similar trends on 8B-scale LLMs). We demonstrate that by generating UTs with UTGen and computing the rate of passing UTs, we can better judge code correctness than using scores from trained, state-of-the-art 8B reward model (RM). For best-of-10 sampling with Qwen2.5 Coder 7B Instruct, UTGen improves accuracy by 4.43% over the RM on HumanEval+. Finally, in Sec. 5.5, we highlight the importance of high-quality unit tests generated by UTGen when debugging with a frontier LLM like GPT-4o (Hurst et al., 2024), finding that using UTGen-feedback from a smaller model, such as Qwen2.5 32B, significantly outperforms self-generated feedback from GPT-4o by nearly 25% (absolute) on the most challenging subset of MBPP+ problems.
@@ 2 Related Work
Automatic Unit Test Generation. Manually writing unit tests is laborious and often infeasible (Chen et al., 2023a; Liu et al., 2024b). Consequently, past research explores automatic UT generation (King, 1976; Cadar et al., 2008; Holler et al., 2012; Cha et al., 2015, inter alia). The advent of LLMs has spurred recent efforts in using them for UT generation (Chen et al., 2023a; Schäfer et al., 2023; Liu et al., 2024b). Specifically, Schäfer et al. (2023) and Liu et al. (2024b) focus on generating unit test inputs via prompting frontier LLMs like GPT-4 and/or iterative prompting, assuming access to the gold solution. In contrast, our models, trained with UTGen, generate both input-output UT pairs based on the task description without relying on the gold implementation. While Chen et al. (2023a) also generate input-output UT pairs using standard LLM prompting, their primary focus is code generation – not the quality of generated UTs. In contrast, we directly model the desiderata (or quality) of UTs, and demonstrate its utility on code generation and debugging.
LLM Debugging. Using LLMs for debugging faulty code, or program repair, has been extensively studied. Debugging approaches are divided into those training models to debug (Moon et al., 2023; Ni et al., 2024; Chae et al., 2024) and those providing external feedback to pretrained models (Chen et al., 2023b; Zhang et al., 2023; Olausson et al., 2023; Zhong et al., 2024). Both rely on gold unit tests for training or feedback. Thus, UTGen complements both methods by providing generated unit tests when human-written tests are scarce or unavailable. In Sec. 3.3, we introduce UTDebug, a debugging pipeline that addresses noisy feedback from inaccurate unit tests through test-time scaling and backtracking. Moreover, in Sec. 5 we show that UTGen’s unit tests can effectively provide feedback to LLMs for code generation and debugging.
3 Unit Test Generation and Automated Debugging
@@ 3.1 Background and Task Setup
Given a natural language task description for a coding task, we focus on generating unit tests (UTs) in the format of input-output pairs that are consistent with the task description . Our setup is consistent with Chen et al. (2023a) and Jain et al. (2024) who also consider unit tests in the form of input-output pairs generated without utilizing the correct implementation of the function. More generally, our setting is akin to parameterized unit testing (Tillmann et al., 2010) and uses the notion of functional correctness, i.e., measuring correctness by simulating an exhaustive set of scenarios (UT inputs) and ensuring that the function performs as expected (as per the problem description ).
@@
3.2 UTGen: Training LLMs for Unit Test Generation
While prior work focuses on curating training data for improving code generation (Guo et al., 2024a; Muennighoff et al., 2023), there is a general lack of dedicated datasets for training the desired UT generator outlined above. Therefore, to improve along the lines of RQ2, we design, UTGen, a training recipe that bootstraps this data from training datasets for code generation i.e., a collection of problem descriptions and their corresponding code.
@@ 3.2 UTGen: Training LLMs for Unit Test Generation
We start with a collection of coding problems with problem descriptions () and gold codes () as illustrated in Fig. 2 (I). Specifically, we use publicly available data from Tulu-3 (Lambert et al., 2024a) due to its large scale and the improvements noted by Lambert et al. (2024a) when finetuning LLMs on coding tasks.222The code solutions in the Tulu-3 SFT mix have either been written by humans or by frontier models and are thus highly likely to be correct (Lambert et al., 2024a). We filter it to focus on Python code with functional abstractions (further details in Appendix B). However, in order to train a unit test generator we would need access to incorrect or faulty code that can be debugged, as unit tests must have some error to attack. We obtain these by using the LLM to perturb the reference code solution.
@@
As mentioned in Sec. 3.1, one of the goals of the unit test generator is to be able to not only generate valid unit tests, but failing unit tests (that trigger errors). To facilitate this, given a problem description, reference code , and buggy candidate code , we sample different unit test inputs (via the prompts in Appendix H). A unit test is failing if , i.e., if the output of the candidate fails to match the reference output (cf. Fig. 2 (II)). Note that, while the LLM can be used to generate the output of unit test, it can often be inaccurate (Jain et al., 2024; Gu et al., 2024); so to ensure output accuracy, we use the output of the reference code during training. Due to the formalization of UT output prediction as a reasoning word problem, we further explore generating chain-of-thought (CoT; Wei et al., 2022) reasoning before predicting the UT output. To this end, we employ the post-hoc rationalization procedure outlined in Zelikman et al. (2022) – given the entire UT , we ask the LLM to generate rationales supporting why is the output corresponding to the input . Then, to create the supervision data, we add these rationales as CoTs prior to the output prediction, as illustrated in Fig. 2 (III). We revisit the discussion of how accurate LLMs are at output prediction during inference in Sec. 5.
Supervised Finetuning.
When training LLMs, the input to the LLM is the same prompt used for sampling unit tests (listed in Appendix H) and the output is a failing unit test. We collect nearly 30K training instances for 8B-scale models and roughly 70K instances for training 32B LLMs (details in Appendix B). The goal here is to improve both output accuracy and attack rate jointly via supervised finetuning using the negative log-likelihood loss using basic hyperparameters outlined in Appendix B.
3.3 UTDebug: Generated Unit Tests for Debugging
A key difference when using generated unit tests, as opposed to human-generated UTs, is the degree of noise in the feedback. Despite training models via UTGen or other methods, a generated unit test may not be 100% accurate. This can manifest in two ways: 1) the generated UT input does not fail for the code under debugging , i.e., ; 2) the generated UT output is inaccurate, i.e., not consistent with what a gold solution would yield (). Both types of errors can negatively impact the utility of unit tests for debugging, as shown in Fig. 3 (left). A non-adversarial input might result in faulty code being misclassified as correct and prematurely removed from debugging. Additionally, incorrect outputs can cause false positives, introducing errors to otherwise correct code. Even if the candidate code has bugs, incorrect UT outputs can lead to incorrect feedback, thereby, degrading performance. These issues motivate our RQ3 to incorporate noisy feedback. We propose UTDebug, which includes two ways to mitigate noisy feedback from automatically generated unit tests (additional details in Appendix C). We empirically validate the importance of both these components of UTDebug in Appendix E.